# Scrutinize What We Ignore: Reining In Task Representation Shift Of Context-Based Offline Meta Reinforcement Learning

### v3

@@ Abstract
Offline meta reinforcement learning (OMRL) has emerged as a promising approach for interaction avoidance and strong generalization performance by leveraging pre-collected data and meta-learning techniques. Previous context-based approaches predominantly rely on the intuition that alternating optimization between the context encoder and the policy can lead to performance improvements, as long as the context encoder follows the principle of maximizing the mutual information between the task variable $M$ and its latent representation $Z$ ($I(Z;M)$) while the policy adopts the standard offline reinforcement learning (RL) algorithms conditioning on the learned task representation. Despite promising results, the theoretical justification of performance improvements for such intuition remains underexplored. Inspired by the return discrepancy scheme in the model-based RL field, we find that the previous optimization framework can be linked with the general RL objective of maximizing the expected return, thereby explaining performance improvements. Furthermore, after scrutinizing this optimization framework, we find it ignores the variation of the task representation in the alternating optimization process, which weakens the condition necessary for monotonic performance improvements, and may therefore violate the monotonicity. We name this issue task representation shift and theoretically prove that the monotonic performance improvements can be guaranteed with appropriate context encoder updates. We use different settings to rein in the task representation shift on three widely adopted training objectives concerning maximizing $I(Z;M)$ across different data qualities. Empirical results show that reining in the task representation shift can indeed improve performance. Our work opens up a new avenue for OMRL, leading to a better understanding between the task representation and performance improvements.
@@ 1 Introduction
RL has driven impressive advances in many complex decision-making problems in recent years [silver2018general, schrittwieser2020mastering, zhang2023safe], primarily through online RL methods. However, the extensive interactions required by online RL entail high costs and safety concerns, posing significant challenges for real-world applications. Offline RL [wu2019behavior, levine2020offline] offers an appealing alternative by efficiently leveraging pre-collected data for policy learning, thereby circumventing the need for online interaction with the environment. This advantage extends the application of RL, covering healthcare [fatemi2022semi, tang2022leveraging], robotics [sinha2022s4rl, pmlr-v164-kumar22a] and games [schrittwieser2021online, mathieu2023alphastar]. Though demonstrating its superiority, offline RL holds a notable drawback towards generalizing to the unknown [ghosh2021generalization].
@@
As a remedy, by combining the meta-learning techniques, OMRL [li2020focal, mitchell2021offline, xu2022prompting] has emerged as an effective training scheme toward strong generalization performance and fast adaptation capability while maintaining the merits of offline RL. Among the OMRL research, context-based OMRL (COMRL) algorithms [li2020focal, li2024towards] hold a popular paradigm that seeks optimal meta-policy conditioning on the context of Markov Decision Processes (MDPs). Specifically, these methods [li2020focal, gao2024context, li2024towards] propose to train a context encoder via maximizing $I(Z;M)$ to learn the task representation from the collected context (e.g. trajectories) and train the downstream policy with standard offline RL algorithms conditioning on the task representation, as shown in Figure 1 left. The context encoder and policy are trained in an alternating manner, with each being updated once per cycle. Despite promising results, the theoretical justification of performance improvements for such intuition remains underexplored.
@@
Figure 1: Our training framework compared to the previous training framework. They both adopt the alternating optimization framework to train the context encoder and the policy. However, ours considers the variation of task representation ignored by the previous training framework by imposing the extra condition to determine whether needs to update the context encoder.
@@
Intriguingly, we formalize this training framework as an alternating two-stage optimization framework and then link it with the general RL objective of maximizing the expected return. To be specific, maximizing $I(Z;M)$ and adopting standard offline RL algorithms can be interpreted as consistently raising the lower bound of the expected return conditioning on the optimal task representation distribution among the expectation of tasks (See Section 4.1). This is achieved by extending the return discrepancy scheme [janner2019trust] to the COMRL framework. Thus, this analysis provides a feasible explanation for the performance improvement guarantee.
@@
More importantly, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation in the alternating process. This would cause the optimization framework to incorrectly conclude that the monotonic performance improvement can be guaranteed with a better approximation to the optimal task representation distribution. By explicitly modeling the variation of task representation, we prove that it is a critical part of monotonic performance improvement. Without this, the monotonicity can be violated. To highlight the characteristic, we name this issue task representation shift and theoretically prove that it is possible to achieve monotonic performance improvement with appropriate updates (See Section 4.2).
@@
To show the impacts of this issue, we use different settings to rein in the task representation shift on contrastive objective adopted in [li2020focal, gao2024context, li2024towards] and reconstruction objective adopted in [zintgraf2019varibad, dorfman2021offline, li2024towards] across different data qualities. These two objectives are proven to maximize the approximate bounds of $I(Z;M)$ [li2024towards]. We also introduce the cross-entropy objective, which is a straightforward discrete approximation to $I(Z;M)$. Empirical results show that reining in the task representation shift can indeed improve performance. Our work opens up a new avenue for OMRL, leading to a better understanding between the task representation and performance improvements.
@@ 2 Related Works
Offline RL. Offline RL [fujimoto2019off] leverages a fully offline dataset to perform policy optimization and value function learning, avoiding explicit interaction with the environment. To overcome the notorious distribution shift and the induced value overestimation issues, prior works adopt methods based on policy regularization [fujimoto2019off, kumar2019stabilizing, fujimoto2021minimalist], value penalty [wu2019behavior, kumar2020conservative] and sequence modeling [chen2021decision, janner2021offline]. While these methods have demonstrated promising performance on a single task, their generalization performance to the unknown remains limited [ghosh2021generalization]. Our work follows the offline RL setting, attempting to adopt meta-learning techniques to improve the generalization performance.
Meta RL. Meta RL has emerged as an enticing avenue for effectively adapting to unknown environments [finn2017model]. Unlike traditional gradient-based meta RL methods [finn2017model, mitchell2021offline], context-based meta RL methods demonstrate superior advantages in both sample efficiency and asymptotic performance [rakelly2019efficient, zintgraf2019varibad, li2024towards]. Our work delves into this promising domain to advance the capabilities of context-based meta RL further.
Context-based Offline Meta RL. As the marriage between context-based meta RL and offline RL, COMRL combines the merits of both sides. Specifically, COMRL methods [li2020focal, yuan2022robust, gao2024context, li2024towards] leverage the offline dataset to train a context encoder to learn robust task representations, and then pass the representation to the policy and value function as input. At test time, COMRL methods leverage the generalization ability of the learned context encoder to perform meta-adaptation. Acknowledging the theoretical insights from [li2024towards], prior COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] mainly focus on how to design a context encoder learning algorithm to better approximate $I(Z;M)$. However, our work moves the focus on how to refine the condition of monotonic performance improvements based on a given context encoder learning algorithm. Center around this motivation, we identify a new issue called task representation shift, which is ignored by the previous COMRL endeavors.
Performance Improvement Guarantee.  Ensuring performance improvement is a key concern in both online and offline reinforcement learning settings. In online RL, performance improvement guarantees are often established through methods such as performance difference bounds [schulman2015trust, kakade2002approximately, ji2022update, zhang2023how], return discrepancy [janner2019trust, luo2018algorithmic], and regret bounds [osband2014model, curi2020efficient]. In offline RL, CQL-based methods [kumar2020conservative, yu2021combo] also enjoy safe policy improvement guarantees. However, most works focus on the single-task setting, leaving the performance improvement guarantees in the COMRL settings largely unexplored. While ContrBAR [choshen2023contrabar] also benefits from the performance improvement guarantee in the OMRL setting, it is specifically served for its own framework. In contrast, our work addresses a broader class of algorithms that maximize $I(Z;M)$, including but not limited to ContrBAR.
@@ 3.1 Problem Statement
A task in RL is generally formalized as a fully observed MDP [puterman2014markov], which is represented by a tuple $(\mathcal{S},\mathcal{A},P,\rho_{0},R,\gamma)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $P(s^{\prime}|s,a)$, reward function $R(s,a)$, initial state distribution $\rho_{0}(s)$ and discount factor $\gamma\in[0,1]$. For arbitrary policy $\pi(a|s)$, we denote the state-action distribution at timestep $t$ as $d_{\pi,t}(s,a)\triangleq\text{Pr}(s_{t}=s,a_{t}=a|s_{0}\sim\rho_{0},a_{t}\sim\pi,s_{t+1}\sim P,\forall t\geq 0)$. The discounted state-action distribution of given $\pi(a|s)$ is denoted as $d_{\pi}(s,a)\triangleq(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}d_{\pi,t}(s,a)$. The ultimate goal is to find an optimal policy $\pi(a|s)$ to maximize the expected return $\mathbb{E}_{(s,a)\sim d_{\pi}}[R(s,a)]$.
Given a distribution of tasks $p(M)$, every task $m$ sampled from $p(M)$ is an MDP sharing the same state space and action space but differing from transition dynamics and reward functions. It is widely assumed that the tasks in the meta RL setting are randomly sampled from $p(M)$ [li2020focal, yuan2022robust, gao2024context, li2024towards]. We denote the number of training tasks as $N$. For each task $i\in[0,1,...,N-1]$, an offline dataset $D_{i}=\{(s_{i,j},a_{i,j},s^{\prime}_{i,j},r_{i,j})\}_{j=1}^{K}$ is collected in advance, where $K$ denotes the number of transitions. The learning algorithm is required to train a meta-policy $\pi_{\text{meta}}$ with access only to the given offline datasets. At test time, given an unseen task, the meta policy $\pi_{\text{meta}}$ performs task adaptation to get a task-specific policy and then will be evaluated in the environment.
@@ 3.2 Context-based OMRL
Previous COMRL endeavors hold that the variation over different tasks can be represented by a compact representation $z$ [yuan2022robust]. In practice, they choose to train a context encoder $Z(\cdot|x)$ to extract the task information [li2020focal, dorfman2021offline], where $x$ denotes the given context. Then, the policy is learned by conditioning on the task representation as $\pi(a|s,Z(\cdot|x);\theta)$. Naturally, there exists an optimal task representation distribution $Z^{*}(\cdot|x)$ that meets the optimal expected return for any policy parameter $\theta$ is determined by $\pi(a|s,Z^{*}(\cdot|x);\theta)$. Hence, the objective of COMRL is formalized as:
$\max_{\theta}J^{*}(\theta)=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]$ (1)
@@
Previous COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] are proven to optimize the context encoder by maximizing the approximate bounds of $I(Z;M)$, as shown in Theorem 3.1. They hold the intuition that using such kind of approach to optimize the context encoder and adopting the standard offline RL algorithms to optimize the policy can lead to performance improvements. However, the theoretical justification of performance improvement for such intuition has been less explored. For later use, we denote $Z_{mutual}(\cdot|x)$ as the optimal solution for maximizing $I(Z;M)$. Whether $Z_{mutual}(\cdot|x)$ is equivalent to $Z^{*}(\cdot|x)$ is still an open problem. Hence, we introduce an extra notation here to avoid confusion.
Theorem 3.1 ([li2024towards]).
@@
where 1) $\mathcal{L}_{FOCAL}\equiv-I(Z;X)=-I(Z;X_{t}|X_{b})-I(Z;X_{b})$; 2) $\mathcal{L}_{CORRO}\equiv-I(Z;X_{t}|X_{b})$; 3) $\mathcal{L}_{CSRO}\geq(\lambda-1)I(Z;X)-\lambda I(Z;X_{t}|X_{b})$
@@ 4 Methods
We first introduce a Lipschitz assumption, a widely used technique in both model-free [song2019efficient, ghosh2022provably] and model-based [ji2022update, zhang2023how] RL frameworks.
@@
$\forall Z_{1}(\cdot|x),Z_{2}(\cdot|x)\in\mathbb{Z},\,\,|\pi(\cdot|s,Z_{1}(\cdot|x);\theta)-\pi(\cdot|s,Z_{2}(\cdot|x);\theta)|\leq L_{z}\cdot|Z_{1}(\cdot|x)-Z_{2}(\cdot|x)|$ (3)
@@ 4.1 A Performance Improvement Perspective Towards Prior Works
Our goal is to arrive at an optimization objective for COMRL to get the performance improvement guarantee. Motivated by the return discrepancy scheme [janner2019trust] in the model-based RL field, we can similarly build a tractable lower bound for $J^{*}(\theta)$.
@@
where $J(\theta)$ is the expected return of the policy $\pi$ among the expectation of tasks conditioning on the learned task embedding $Z(\cdot|x)$ as $J(\theta)=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]]$.
From Definition 4.2, we know that if the estimation error $|J^{*}(\theta)-J(\theta)|$ can be upper-bounded, the lower bound for $J^{*}(\theta)$ can be established simultaneously. Then, Eq. (4) can be interpreted as a unified training objective for both the context encoder and the policy. Specifically, we can alternate the optimization target to lift the lower bound for $J^{*}(\theta)$, where the holistic optimization process can be seen as a two-stage alternating framework with the first stage updating the context encoder to minimize $|J^{*}(\theta)-J(\theta)|$ and the second stage updating the policy to maximize $J(\theta)-|J^{*}(\theta)-J(\theta)|$ as $J^{*}(\theta)\geq J(\theta)-|J^{*}(\theta)-J(\theta)|$. Hence, the performance improvement guarantee can be achieved.
The following theorem induces a tractable upper bound for $|J^{*}(\theta)-J(\theta)|$, connecting the bridge between the intuition of previous works and the theoretical justification concerning the performance improvement guarantee.
@@
$|J^{*}(\theta)-J(\theta)|\leq\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (5)
Proof: See Section 8.2 in the Appendix. ∎
@@
$J^{*}(\theta)\geq J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (6)
Note that $|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|$ is a constant w.r.t the learned task representation $Z(\cdot|x)$. Hence, this term can be ignored when optimizing $Z(\cdot|x)$. Admittedly, there may exist a smarter algorithm to further reduce the gap of $|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|$, but this is not our main focus and we leave this to future work.
Theorem 4.3 and Eq. (6) indicate the following two principles for the context encoder and the policy. For the context encoder learning, we should minimize the return bound, namely minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$. For the policy learning, we should lift the lower bound of $J^{*}(\theta)$, namely maximizing $J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)\equiv J(\theta)$, as the task representation related terms are constants for policy optimization. To maximize $J(\theta)$, we can adopt standard offline RL algorithms [wu2019behavior, fujimoto2021minimalist, kumar2020conservative] similar to [yang2022unified].
Recall that previous COMRL works adopt alternatively training the context encoder by maximizing $I(Z;M)$, which can be approximately seen as minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$ and training the policy by conditioning on the learned task representation as well as applying the standard offline RL algorithms. Therefore, we argue that previous COMRL works can be interpreted by this framework, thus providing an explanation concerning performance improvement guarantee.
However, this optimization framework only considers the discrepancy between $Z(\cdot|x)$ and $Z_{mutual}(\cdot|x)$, without considering the variation of $Z(\cdot|x)$. In the next section, we will show that this may violate the monotonicity of the performance improvements.

### v4

@@ Abstract
Offline meta reinforcement learning (OMRL) has emerged as a promising approach for interaction avoidance and strong generalization performance by leveraging pre-collected data and meta-learning techniques. Previous context-based approaches predominantly rely on the intuition that alternating optimization between the context encoder and the policy can lead to performance improvements, as long as the context encoder follows the principle of maximizing the mutual information between the task variable $M$ and its latent representation $Z$ ($I(Z;M)$) while the policy adopts the standard offline reinforcement learning (RL) algorithms conditioning on the learned task representation. Despite promising results, the theoretical justification of performance improvements for such intuition remains underexplored. Inspired by the return discrepancy scheme in the model-based RL field, we find that the previous optimization framework can be linked with the general RL objective of maximizing the expected return, thereby explaining performance improvements. Furthermore, after scrutinizing this optimization framework, we observe that the condition for monotonic performance improvements does not consider the variation of the task representation. When these variations are considered, the previously established condition may no longer be sufficient to ensure monotonicity, thereby impairing the optimization process. We name this issue task representation shift and theoretically prove that the monotonic performance improvements can be guaranteed with appropriate context encoder updates. We use different settings to rein in the task representation shift on three widely adopted training objectives concerning maximizing $I(Z;M)$ across different data qualities. Empirical results show that reining in the task representation shift can indeed improve performance. Our work opens up a new avenue for OMRL, leading to a better understanding between the task representation and performance improvements.  [^1] Code: https://github.com/betray12138/Task-Representation-Shift
@@ 1 Introduction
RL has driven impressive advances in many complex decision-making problems in recent years (Silver et al., 2018; Schrittwieser et al., 2020; Zhang et al., 2023b; 2024), primarily through online RL methods. However, the extensive interactions required by online RL entail high costs and safety concerns, posing significant challenges for real-world applications. Offline RL (Wu et al., 2019; Levine et al., 2020) offers an appealing alternative by efficiently leveraging pre-collected data for policy learning, thereby circumventing the need for online interaction with the environment. This advantage extends the application of RL, covering healthcare (Fatemi et al., 2022; Tang et al., 2022), robotics (Sinha et al., 2022; Kumar et al., 2022) and games (Schrittwieser et al., 2021). Though demonstrating its superiority, offline RL holds a notable drawback towards generalizing to the unknown (Ghosh et al., 2021).
@@
As a remedy, by combining the meta-learning techniques, OMRL (Li et al., 2020; Mitchell et al., 2021; Xu et al., 2022) has emerged as an effective training scheme toward strong generalization performance and fast adaptation capability while maintaining the merits of offline RL. Among the OMRL research, context-based OMRL (COMRL) algorithms (Li et al., 2020; 2024) hold a popular paradigm that seeks optimal meta-policy conditioning on the context of Markov Decision Processes (MDPs). Specifically, these methods (Li et al., 2020; Gao et al., 2024; Li et al., 2024) propose to train a context encoder via maximizing $I(Z;M)$ to learn the task representation from the collected context (e.g. trajectories) and train the downstream policy with standard offline RL algorithms conditioning on the task representation, as shown in Figure 1 left. The context encoder and policy are trained in an alternating manner, with each being updated once per cycle. Despite promising results, the theoretical justification of performance improvements for such intuition remains underexplored.
@@
Figure 1: Our training framework compared to the previous training framework. They both adopt the alternating optimization framework to train the context encoder and the policy. However, our training framework considers the previously ignored variation of task representation by introducing an extra condition to decide whether the context encoder should be updated.
@@
Intriguingly, we formalize this training framework as an alternating two-stage optimization framework and then link it with the general RL objective of maximizing the expected return. To be specific, maximizing $I(Z;M)$ and adopting standard offline RL algorithms can be interpreted as consistently raising the lower bound of the expected return conditioning on the optimal task representation distribution (See Section 4.1). This is achieved by extending the return discrepancy scheme (Janner et al., 2019) to the COMRL framework. Thus, this analysis provides a feasible explanation for the performance improvement guarantee.
@@
More importantly, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation in the alternating process. This would cause the optimization framework to incorrectly conclude that the monotonic performance improvement can be guaranteed with a better approximation to the optimal task representation distribution. By explicitly modeling the variation of task representation, we conclude that it is a critical part of monotonic performance improvement. To highlight the characteristic, we name this issue task representation shift and theoretically prove that it is possible to achieve monotonic performance improvement with appropriate context encoder updates (See Section 4.2).
@@
To show the impacts of this issue, we use different settings to rein in the task representation shift on three widely adopted objectives concerning maximizing $I(Z;M)$, covering the upper bound, the lower bound, and the direct approximation. Empirical results show that reining in the task representation shift can indeed improve performance. Our work opens up a new avenue for OMRL, leading to a better understanding between the task representation and performance improvements.
@@ 2 Related Works
Context-Based Offline Meta RL. As the marriage between context-based meta RL and offline RL, COMRL combines the merits of both sides. Specifically, COMRL methods (Li et al., 2020; Yuan & Lu, 2022; Gao et al., 2024; Li et al., 2024) leverage the offline dataset to train a context encoder to learn robust task representations, and then pass the representation to the policy and value function as input. At test time, COMRL methods leverage the generalization ability of the learned context encoder to perform meta-adaptation. According to the theoretical insights from (Li et al., 2024), prior COMRL works (Li et al., 2020; Yuan & Lu, 2022; Gao et al., 2024; Li et al., 2024) mainly focus on designing a context encoder learning algorithm to better approximate $I(Z;M)$. However, our work turns the focus to refining the condition of monotonic performance improvements based on a given context encoder learning algorithm. Centering around this motivation, we identify a new issue called task representation shift, which is ignored by the previous COMRL endeavors.
Performance Improvement Guarantee.  Ensuring performance improvement is a key concern in both online and offline reinforcement learning settings. In online RL, performance improvement guarantees are often established through methods such as performance difference bounds (Kakade & Langford, 2002; Schulman et al., 2015; Ji et al., 2022; Zhang et al., 2023a), return discrepancy (Luo et al., 2018; Janner et al., 2019), and regret bounds (Osband & Van Roy, 2014; Curi et al., 2020). In offline RL, CQL-based methods (Kumar et al., 2020; Yu et al., 2021) also enjoy safe policy improvement guarantees. However, most works focus on the single-task setting, leaving the performance improvement guarantees in the context-based meta-RL settings largely unexplored. While ContrBAR (Choshen & Tamar, 2023) also benefits from the performance improvement guarantee, the theoretical insight focuses on the online setting. Additionally, it is tailored to one particular approximation of $I(Z;M)$ as it makes assumptions specific to this approximation. In contrast, our work focuses on the general offline setting, addressing a broader class of algorithms that maximize various bounds of $I(Z;M)$.
@@ 3.1 Problem Statement
A task in RL is generally formalized as a fully observed MDP (Puterman, 2014), which is represented by a tuple $(\mathcal{S},\mathcal{A},P,\rho_{0},R,\gamma)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $P(s^{\prime}|s,a)$, reward function $R(s,a)$, initial state distribution $\rho_{0}(s)$ and discount factor $\gamma\in[0,1]$. For arbitrary policy $\pi(a|s)$, we denote the state-action distribution at timestep $t$ as $d_{\pi,t}(s,a)\triangleq\text{Pr}(s_{t}=s,a_{t}=a|s_{0}\sim\rho_{0},a_{t}\sim\pi,s_{t+1}\sim P,\forall t\geq 0)$. The discounted state-action distribution of given $\pi(a|s)$ is denoted as $d_{\pi}(s,a)\triangleq(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}d_{\pi,t}(s,a)$. The ultimate goal is to find an optimal policy $\pi(a|s)$ to maximize the expected return $\mathbb{E}_{(s,a)\sim d_{\pi}}[R(s,a)]$.
It is widely assumed that the task $m$ in the meta RL setting is randomly sampled from the task distribution $p(M)$ (Li et al., 2020; Yuan & Lu, 2022; Gao et al., 2024; Li et al., 2024). We denote the number of training tasks as $N$. For each task $i\in[0,1,...,N-1]$, an offline dataset $D_{i}=\{(s_{i,j},a_{i,j},s^{\prime}_{i,j},r_{i,j})\}_{j=1}^{K}$ is collected in advance, where $K$ denotes the number of transitions. The learning algorithm is required to train a meta-policy $\pi_{\text{meta}}$ with access only to the given offline datasets. At test time, given an unseen task, the meta policy $\pi_{\text{meta}}$ performs task adaptation to get a task-specific policy and then will be evaluated in the environment.
@@ 3.2 Context-based OMRL
Previous COMRL endeavors employ a compact representation $z$ to capture/quantify the variation over tasks (Yuan & Lu, 2022). In practice, they choose to train a context encoder $Z(\cdot|x;\phi)$ to extract the task information (Li et al., 2020; Dorfman et al., 2021), where $x$ denotes the given context and $\phi$ denotes the parameters of the context encoder. Then, the policy is learned by conditioning on the task representation as $\pi(a|s,Z(\cdot|x;\phi);\theta)$. We assume there exists an optimal task representation distribution $Z(\cdot|x;\phi^{*})$ that meets the optimal expected return for any policy parameter $\theta$ is determined by $\pi(a|s,Z(\cdot|x;\phi^{*});\theta)$. Hence, the objective of COMRL is formalized as:
$\max_{\theta}J^{*}(\theta)=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x;\phi^{*});\theta)}}[R_{m}(s,a)]]$ (1)
@@
Previous COMRL works (Li et al., 2020; 2021; Yuan & Lu, 2022; Gao et al., 2024; Li et al., 2024) are proven to optimize the context encoder by maximizing the approximate bounds of $I(Z;M)$, as shown in Theorem 3.1. They hold the intuition that using such kind of approach to optimize the context encoder and adopting the standard offline RL algorithms to optimize the policy can lead to performance improvements. However, the theoretical justification of performance improvement for such intuition has been less explored. For later use, we denote $Z(\cdot|x;\phi^{mutual})$ as the optimal solution for maximizing $I(Z;M)$. Whether $Z(\cdot|x;\phi^{mutual})$ is equivalent to $Z(\cdot|x;\phi^{*})$ remains an open problem. To avoid confusion, we introduce an extra notation here.
Theorem 3.1 ((Li et al., 2024)).
@@
where 1) $\mathcal{L}_{FOCAL}\equiv-I(Z;X)=-I(Z;X_{t}|X_{b})-I(Z;X_{b})$; 2) $\mathcal{L}_{CORRO}\equiv-I(Z;X_{t}|X_{b})$; 3) $\mathcal{L}_{CSRO}\geq(\lambda-1)I(Z;X)-\lambda I(Z;X_{t}|X_{b})$, and $\equiv$ denotes equality up to a constant.
@@ 4 Methods
We first introduce a Lipschitz assumption, a widely used technique in both model-free (Song & Sun, 2019; Ghosh et al., 2022) and model-based (Ji et al., 2022; Zhang et al., 2023a) RL frameworks.
@@
$\forall Z(\cdot|x;\phi_{1}),Z(\cdot|x;\phi_{2})\in\mathbb{Z},\,\,|\pi(\cdot|s,Z(\cdot|x;\phi_{1});\theta)-\pi(\cdot|s,Z(\cdot|x;\phi_{2});\theta)|\leq L_{z}\cdot|Z(\cdot|x;\phi_{1})-Z(\cdot|x;\phi_{2})|$ (3)
The proofs for all the following sections are provided in Appendix Section 8.2.
@@ 4.1 A Performance Improvement Perspective Towards Prior Works
Our goal is to arrive at an optimization objective for COMRL to get the performance improvement guarantee. Motivated by the return discrepancy scheme (Janner et al., 2019) in the model-based RL field, we can similarly build a tractable lower bound for $J^{*}(\theta)$.
@@
where $J(\theta)=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x;\phi);\theta)}}[R_{m}(s,a)]]$ is the expected return of the policy $\pi$ conditioning on the learned task representation $Z(\cdot|x;\phi)$.
From Definition 4.2, we know that if the estimation error $|J^{*}(\theta)-J(\theta)|$ can be upper-bounded, the lower bound for $J^{*}(\theta)$ can be established simultaneously. Then, Eq. (4) can be interpreted as a unified training objective for both the context encoder and the policy. Specifically, we can alternate the optimization target to lift the lower bound for $J^{*}(\theta)$, where the whole optimization process can be seen as a two-stage alternating framework with the first stage updating the context encoder to minimize $|J^{*}(\theta)-J(\theta)|$ and the second stage updating the policy to maximize $J(\theta)-|J^{*}(\theta)-J(\theta)|$ as $J^{*}(\theta)\geq J(\theta)-|J^{*}(\theta)-J(\theta)|$. Hence, the performance improvement guarantee can be achieved.
The following theorem induces a tractable upper bound for $|J^{*}(\theta)-J(\theta)|$, bridging the gap between the intuition of previous works and the theoretical justification concerning the performance improvement guarantee.
@@
$|J^{*}(\theta)-J(\theta)|\leq\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x;\phi)-Z(\cdot|x;\phi^{mutual})|+|Z(\cdot|x;\phi^{mutual})-Z(\cdot|x;\phi^{*})|)$ (5)
@@
$J^{*}(\theta)\geq J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x;\phi)-Z(\cdot|x;\phi^{mutual})|+|Z(\cdot|x;\phi^{mutual})-Z(\cdot|x;\phi^{*})|)$ (6)
Note that $|Z(\cdot|x;\phi^{mutual})-Z(\cdot|x;\phi^{*})|$ is a constant w.r.t the learned task representation $Z(\cdot|x;\phi)$. Hence, this term can be ignored when optimizing $Z(\cdot|x;\phi)$. Admittedly, there may exist a smarter algorithm to further reduce the gap of $|Z(\cdot|x;\phi^{mutual})-Z(\cdot|x;\phi^{*})|$, but this is not our main focus and we leave this to future work.
Theorem 4.3 and Eq. (6) indicate the following two principles for the context encoder and the policy. For the context encoder learning, we should minimize the return bound, namely minimizing $\mathbb{E}_{m,x}|Z(\cdot|x;\phi)-Z(\cdot|x;\phi^{mutual})|$. For the policy learning, we should lift the lower bound of $J^{*}(\theta)$, namely maximizing $J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x;\phi)-Z(\cdot|x;\phi^{mutual})|+|Z(\cdot|x;\phi^{mutual})-Z(\cdot|x;\phi^{*})|)\equiv J(\theta)$, as the task representation related terms are constants for policy optimization. To maximize $J(\theta)$, we can adopt standard offline RL algorithms (Wu et al., 2019; Kumar et al., 2020; Fujimoto & Gu, 2021) similar to (Yang et al., 2022).
Recall that previous COMRL works adopt alternatingly training the context encoder by maximizing $I(Z;M)$, which can be approximately seen as minimizing $\mathbb{E}_{m,x}|Z(\cdot|x;\phi)-Z(\cdot|x;\phi^{mutual})|$ and training the policy by conditioning on the learned task representation as well as applying the standard offline RL algorithms. Therefore, we argue that our theoretical analyses provide an explanation for the performance improvement guarantee of previous COMRL works.
However, this optimization framework only considers the discrepancy between $Z(\cdot|x;\phi)$ and $Z(\cdot|x;\phi^{mutual})$, without considering the variation of $Z(\cdot|x;\phi)$. In the next section, we will show that this may violate the monotonicity of the performance improvements.